ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 252, 2022

Bainbridge, M., C. Johnson, C. Judge and I. Park Haupt’s theorem for strata of abelian differentials

429

Benhamou, T. and M. Gitik Intermediate models of Magidor– Radin forcing. I

47

Brassil, M. and Z. Reichstein A graph-theoretic approach to a conjecture of Dixon and Pressman

291

Castillo, J. M. F., W. Cuellar, V. Ferenczi and Y. Moreno On Disjointly singular centralizers

215

Cuellar, W. See J. M. F. Castillo Ferenczi, V. See J. M. F. Castillo Gargate, I. G. and T. C. de Mello Images of multilinear

polynomials on n × n upper triangular matrices over infinite fields

337

Gitik, M. See T. Benhamou Johnson, C. See M. Bainbridge Judge, C. See M. Bainbridge

Kazukawa, D. Convergence of metric transformed spaces  243

Kihara, T. Topological reducibilities for discontinuous functions and their structures

461

Kühne, L. and G. Yashfe Representability of matroids by c-arrangements is undecidable

95

Lubetzky, E. and Y. Peled Noise sensitivity of critical random graphs

187

de Mello, T. G. See I. G. Gargate Moreno, Y. See J. M. F. Castillo Park, I. See M. Bainbridge Peled, Y. See E. Lubetzky Phung, X. K. On dynamical finiteness properties of algebraic group

shifts

355

Rasila, A. See Q. Zhou Reichstein, Z. See M. Brassil

Trost, A. A. Strong boundedness of split Chevalley groups  1

Varma, I. The mean number of 3-torsion elements in ray class groups of quadratic fields

149

Yashfe, G. See L. Kühne Zhou, Q. and A. Rasila Teichmüller’s problem for Gromov

hyperbolic domains 399